OpenStudy (anonymous):

Find the area A inside r = 1 + cos θ and outside r = cos θ

OpenStudy (anonymous):

The total area is given by the integral \[\large2\cdot\left[\frac{1}{2}\int_0^{\pi}\left((1+\cos\theta)^2-(\cos\theta)^2\right)\;d\theta\right]\] The 2 comes from the fact that the curves are symmetric across the x-axis.